I'm doing algebra right now and I've looked threw the book for these few questions and it's impossible . Like this one for example 3(x-4)=5(x+2)
Answer by richard1234(5390)

It is definitely not impossible. We can expand both sides using the distributive property:

3(x-4) = 5(x+2)
3x-12 = 5x+10

Subtract 3x from both sides.

-12 = 2x+10

Subtract 10 from both sides.

-22 = 2x

Divide both sides by 2 to get x = -11. You can check by replacing x with -11.

I recommend you reread your algebra book for the examples. Obviously, the examples won't be the same ones as the ones you will have to do, but you will learn how to use these examples to solve new (and eventually challenging) problems.